### v1

@@ 2.4 Limitations in Consistency, Hallucination, and Generalization
Another recurrent issue is inconsistency. Identical or nearly identical inputs can elicit divergent responses, even within the same session. This instability hinders reproducibility in simulated environments and challenges the interpretability of agent behavior. Wu et al. (2025) emphasizes that LLM agents, when evaluated across multiple instantiations, display lower variance and reduced behavioral diversity compared to human participants, while also amplifying social stereotypes.
@@ 2.5 Summary and Implications
3 LLMs in Social Simulation
@@
The use of LLMs in social simulations builds upon years of research that have seen the application of machine learning techniques in multi-agent systems and, by extension, in agent-based models (Drogoul & Zucker 1998). It is therefore essential to understand whether LLMs represent a major breakthrough or, more modestly, a logical evolution of this partnership, and to determine if they share the same methodological challenges as previous approaches or introduce new ones. Building on the previous discussion regarding the representational capacities of LLMs in psychological and behavioral contexts, this section examines their application in the field of social simulation. We focus on how they are currently used to instantiate artificial societies, the architectural strategies adopted for agent based modeling, and the methodological challenges related to validation and scaling.
3.1 Foundational and Emerging Platforms
@@
A foundational illustration of LLM-based social simulation is the Generative Agents framework introduced by Park (2023), widely known as the “Smallville” project. In this system, twenty-five agents inhabit a sandbox environment reminiscent of The Sims, where they autonomously engage in daily routines, form interpersonal relationships, and generate emergent collective behaviors. Each agent maintains a natural language memory stream, allowing it to reflect on past events, summarize experiences, and formulate plans. Remarkably, the simple assignment of a goal—such as organizing a party—was sufficient to initiate a chain of spontaneous, coordinated interactions among agents. Evaluation efforts focused primarily on assessing the believability and internal coherence of these emergent dynamics, rather than on their correspondence to empirical data.
@@
More recent efforts have extended this line of research to address questions of scale, complexity, and empirical grounding. One prominent example is AgentSociety, a large-scale platform designed to simulate the dynamics of human societies with over 10,000 LLM-based agents (Piao et al. 2025). Each agent is instantiated using GPT-4 and equipped with cognitive modules for memory, goal management, decision-making, and social relationship tracking. These agents interact within rich environments encompassing spatial mobility, economic activity, social networks, and exogenous events such as natural disasters or political crises. AgentSociety has been applied to simulate phenomena including political polarization, rumor diffusion, and the societal impacts of universal basic income. Its validation strategy emphasizes empirical fidelity by replicating outcomes observed in behavioral experiments and large-scale social surveys.
@@
Complementing these large-scale efforts, Simulate Anything offers a flexible platform for generating demographically diverse agent populations using LLMs, leveraging real-world user data for initialization and sampling (Wang 2025). The system supports scenario-specific tuning and hypothesis testing across domains such as politics, journalism, and economics. Its principal contribution lies in enabling heterogeneous role configurations and behavioral variability, though its validation methodologies are comparatively underdeveloped.
@@
Another significant contribution in this domain is S3 (Social-network Simulation System), which specifically focuses on social network dynamics through LLM-empowered agents (Gao et al. 2025). S3 demonstrates particular strengths in simulating three pivotal aspects of human social behavior: emotion, attitude, and interaction behaviors. It leverages advanced prompt engineering and prompt tuning techniques to ensure authentic human-like agent behavior within social networks.
@@
Several more recent platforms have further advanced the field in terms of architectural modularity and methodological rigor. GenSim, for instance, is a general-purpose simulation engine that supports tens of thousands of LLM agents (Tang et al. 2025). It introduces structured abstractions for social routines, long-term memory management, and adaptive behavior correction, all orchestrated by a dynamic scheduling system. Validation combines empirical replication, face validity assessments, and anomaly detection to ensure both plausibility and statistical consistency.
@@
Another emerging contribution is AgentTorch, a modular and scalable framework that unifies large-scale agent-based modeling with differentiable programming and LLM-powered behavior synthesis (Chopra et al. 2023). Instead of relying on expensive real-time prompting for each agent, AgentTorch uses a small number of archetypal LLM agents to generate compact, reusable policy representations. These are then deployed across massive populations of lightweight agents in GPU-accelerated simulations. Whereas narrative-first systems emphasize fine-grained agent individuality, AgentTorch focuses on experimental control, statistical robustness, and scientific reproducibility, making it particularly suited for large-scale hypothesis testing and parameter exploration in computational social science.
@@
SALLMA (Scalable Architecture for LLM Multi-Agent Applications) adopts a layered architecture, separating operational processes (intent formation, task execution, communication) from knowledge-level components (agent profiles, shared memory, workflows) (Becattini et al. 2025). This modular design facilitates agent specialization and integration with external tools via APIs. Its validation pipeline combines automated benchmarking with expert-in-the-loop review, aided by interactive dashboards and simulation playback for iterative refinement.
@@
SocioVerse represents a shift toward population-scale calibration (Zhang et al. 2025b). It initializes agents from millions of real-world user profiles to mirror authentic distributions of attitudes, preferences, and social structures. The platform supports fine-grained tuning for scenario exploration, such as testing the effects of policy changes across diverse subpopulations. Its validation strategy emphasizes both empirical benchmarking and participatory evaluation, particularly focusing on the accurate representation of minority groups and on interactive adjustment of simulation parameters.
@@
Similarly, the LLM-AIDSim framework integrates LLMs (Llama3:8b) into ABM for influence diffusion, enabling language-level responses and deeper insights into user agent interactions (Zhang et al. 2025a).
@@
3.2 Architectures for LLM-Based Multi-Agent Systems
3.2.1 Agent Representation and Cognitive Modules
LLM-based multi-agent models generally instantiate each agent as a distinct LLM instance or session using prompt engineering. Architectures typically involve three interconnected modules: a memory system that encodes and retrieves past experiences; a reflection and summarization layer that condenses observations into mental models; and a planning component that uses these reflections to generate actions (Park 2023; Li et al. 2024). These modules are often implemented as chained or nested prompts, with each step leveraging the LLM’s generative capacity to simulate cognitive processes such as recall, reasoning, and intention formation. For instance, the Humanized Agent-Based Models (h-ABM) framework, proposed by Atchade et al. (2024), outlines a modular architecture for integrating LLMs into ABMs to achieve more cognitively and emotionally realistic agents.
Memory Systems and Psychological Foundations
Many LLM-based agent architectures draw inspiration from cognitive psychology, particularly theories of episodic and semantic memory. For example, the memory systems in Generative Agents and AgentSociety are designed to mimic human-like recall, using vector databases or embedding-based retrieval to surface relevant past experiences. Some frameworks explicitly reference psychological models, such as the Atkinson-Shiffrin multi-store model or the concept of “theory of mind,” to inform the design of agent cognition and social reasoning (Li et al. 2024; Doran & et al. 2023). Reflection modules often prompt the LLM to summarize recent events, extract intentions, or infer the mental states of other agents, thus operationalizing constructs from social and developmental psychology.
Established cognitive architectures such as SOAR (Laird 2012) and ACT-R (Anderson et al. 2004) have provided theoretical foundations for agent design. SOAR, developed by John Laird, Allen Newell, and Paul Rosenbloom at Carnegie Mellon University (Laird 2012), embodies the problem space hypothesis, suggesting that all goal-oriented behavior can be characterized as a search in a problem space defined by states, operators, and goals. This cognitive architecture divides knowledge into procedural memory (skills and know-how), semantic memory (facts about the world), and episodic memory (memories of experiences). Similarly, ACT-R (Adaptive Control of Thought-Rational) (Anderson et al. 2004) posits that human cognition can be modeled through interconnected cognitive and perceptual modules that operate according to rational principles. These architectures inform the contemporary design of LLM agents by providing structured frameworks for organizing memory, decision-making, and learning.
Planning and Decision-Making
@@ 3.2.1 Agent Representation and Cognitive Modules
A hybrid approach is exemplified by the LGC-MARL framework proposed by Jia et al. (2025), which combines an LLM-based planner with graph-structured multi-agent reinforcement learning. In this architecture, the LLM is tasked with decomposing high-level instructions into executable subtasks, evaluating their coherence and feasibility, and formulating collaborative strategies among agents before final action selection. In contrast, more conventional social simulation systems rely predominantly on prompt-engineering strategies to define agent roles, goals, and reasoning pathways, typically without integrating reinforcement signals or dynamic adaptation mechanisms (Junprung 2023). These approaches underscore the current design trade-off between architectural complexity and ease of interpretability in the planning components of LLM-driven simulations.
Orchestration and Communication
@@
Approaches to orchestration and communication vary considerably depending on the complexity and goals of the simulation. For example, platforms like Generative Agents and AgentSociety utilize orchestration layers to manage complex interactions among thousands of agents, often based on sequential flows or perception-reflection-action loops. Others, such as GenSim, integrate structured abstractions for social routines and a dynamic scheduling system. More advanced frameworks, like AgentTorch, move away from costly sequential dependencies by generating reusable policy representations from a small number of archetypal LLM agents, which are then deployed at scale on lightweight agents for GPU-accelerated simulations. SALLMA proposes a layered architecture, separating operational processes (intention formation, task execution) from knowledge-level components (agent profiles, shared memory, workflows), thereby facilitating agent specialization and integration with external tools via APIs. These platforms often support modular ”blocks,” which are reusable behavioral units that allow for systematic experimentation and increased extensibility. Integration with external APIs (e.g., web search, databases, calculators) is also crucial, as it enables agents to access additional information and tools, thereby augmenting their reasoning and decision-making capabilities (Chen et al. 2024b).
@@ 3.2.2 Specialization and Behavioral Diversity
LLM-based multi-agent system (MAS) frameworks are progressively integrating sophisticated specialization protocols, allowing individual agents to adopt roles characterized by unique behavioral repertoires. This specialization is typically realized through several mechanisms: the application of tailored prompt templates that guide an agent’s responses (e.g., (Wei et al. 2022)), which can enforce specific communication styles or decision-making heuristics; the implementation of role-specific memory structures that store relevant knowledge and past interactions (e.g., (Park 2023)), enabling agents to leverage historical data pertinent to their designated function; and scenario-driven fine-tuning processes that adapt an agent’s core model to specific operational contexts (e.g., (Ouyang et al. 2022)), thereby optimizing their performance for particular tasks or environments. Furthermore, some cutting-edge systems are exploring methods to inject personality traits or demographic attributes into agents. This can be achieved either by conditioning prompts with specific personality descriptors (e.g., ”act as a skeptical researcher”) or by sampling attributes from real-world survey distributions (e.g., age, occupation, cultural background), thereby enhancing the realism and diversity of agent behaviors and fostering more nuanced social dynamics (Wang et al. 2024; Chen et al. 2024b). These innovative techniques clearly enhance existing population synthesis algorithms (e.g., (Chapuis et al. 2022)) and enable researchers to explore complex emergent behaviors within agent populations, thereby gaining deeper insights into human-like interactions.
3.3 Validation Strategies and Data Sources
@@ 3.3.1 Empirical Validation and Benchmarking
The validation of LLM-based social simulations remains an active and evolving area of research. Among the most rigorous efforts, frameworks like AgentSociety adopt multi-faceted validation strategies that compare simulated outputs with real-world data. Empirical benchmarking typically involves aligning simulation results with social indicators such as voter turnout, economic behavior, or survey-derived attitudinal distributions from established instruments like the General Social Survey or the World Values Survey (Piao et al. 2025). In addition, some studies replicate classical experimental paradigms—such as public goods games or rumor propagation scenarios—within the simulation environment, and evaluate whether the emergent outcomes converge toward established empirical findings.
Longitudinal consistency is also a critical focus, particularly in simulations involving extended time horizons. Researchers analyze whether agent behavior remains stable and coherent across simulation runs, and whether emergent patterns are reproducible under minor changes to initial conditions or prompt phrasing (Wu et al. 2025). These approaches aim to assess not only face plausibility but also the robustness and generalizability of the simulated social dynamics.
@@ 3.3.2 Human-in-the-Loop Evaluation
Calibration against real-world survey data constitutes another widely used strategy. Here, demographic or attitudinal profiles from empirical sources are used to condition LLM agents, and human evaluators subsequently assess the extent to which the simulated agents reflect the intended population characteristics (Chen et al. 2024b). Nevertheless, recent findings suggest that human raters may be influenced by non-epistemic factors—such as assertiveness, linguistic fluency, or the perceived authority of LLM outputs—leading to potential overestimation of their factual or analytical accuracy (Chen et al. 2024a).

### v2

@@ 2.4 Intrinsic Instability: Hallucinations and Inconsistency
Compounding this issue is the problem of stochastic inconsistency. Identical or nearly identical inputs can elicit divergent responses, even within the same session, which hinders reproducibility and challenges the interpretability of agent behavior [23]. This instability can leave lasting impressions of unreliability, limiting the trustworthiness of the simulation [25]. Furthermore, while LLMs demonstrate superior performance in unrealistic, ’omniscient’ settings where all agents possess full information [1], they struggle significantly in scenarios characterizing real-world conditions, such as information asymmetry or the management of private knowledge [86]. Paradoxically, despite this individual instability, Wu et al. (2025) emphasize that LLM agents, when evaluated across multiple instantiations, often display lower variance and reduced behavioral diversity compared to human participants, amplifying social stereotypes and tending towards a ’mode collapse’ [81]
@@ 2.5 Summary and Implications
3 State of the Art: LLM-Based Social Simulation
@@
The use of LLMs in social simulations builds upon years of research that have seen the application of machine learning techniques in multi-agent systems and, by extension, in agent-based models [22]. It is therefore essential to understand whether LLMs represent a major breakthrough or, more modestly, a logical evolution of this partnership, and to determine if they share the same methodological challenges as previous approaches or introduce new ones. Building on the previous discussion regarding the representational capacities of LLMs in psychological and behavioral contexts, this section examines their application in the field of social simulation. We focus on how they are currently used to instantiate artificial societies, the architectural strategies adopted for agent based modeling, and the methodological challenges related to validation and scaling.
3.1 From Generative Agents to Population-Scale Simulation
@@
A foundational illustration of LLM-based social simulation is the Generative Agents framework introduced by [54], widely known as the “Smallville” project. In this system, twenty-five agents inhabit a sandbox environment reminiscent of The Sims, where they autonomously engage in daily routines, form interpersonal relationships, and generate emergent collective behaviors. Each agent maintains a natural language memory stream, allowing it to reflect on past events, summarize experiences, and formulate plans. Remarkably, the simple assignment of a goal—such as organizing a party—was sufficient to initiate a chain of spontaneous, coordinated interactions among agents. Evaluation efforts focused primarily on assessing the believability and internal coherence of these emergent dynamics, rather than on their correspondence to empirical data.
@@
More recent efforts have extended this line of research to address questions of scale, complexity, and empirical grounding. One prominent example is AgentSociety, a large-scale platform designed to simulate the dynamics of human societies with over 10,000 LLM-based agents [56]. Each agent is instantiated using GPT-4 and equipped with cognitive modules for memory, goal management, decision-making, and social relationship tracking. These agents interact within rich environments encompassing spatial mobility, economic activity, social networks, and exogenous events such as natural disasters or political crises. AgentSociety has been applied to simulate phenomena including political polarization, rumor diffusion, and the societal impacts of universal basic income. Its validation strategy emphasizes empirical fidelity by replicating outcomes observed in behavioral experiments and large-scale social surveys.
Beyond general-purpose social interaction, emerging frameworks are now targeting specific socio-technical systems where physical constraints are paramount. Notably, GATSim (Generative Agent Transport Simulation) is a specialized cognitive framework for urban mobility simulation [49]. Moving away from traditional utility-based discrete choice models, their generative agents demonstrate the ability to plan complex, realistic daily travel itineraries. This work highlights a crucial evolution: the application of LLMs is shifting from purely conversational environments (like Smallville) to spatially constrained domains, where agents must balance personal preferences with physical transportation networks .
@@
Complementing these large-scale efforts, Simulate Anything offers a flexible platform for generating demographically diverse agent populations using LLMs, leveraging real-world user data for initialization and sampling [78]. The system supports scenario-specific tuning and hypothesis testing across domains such as politics, journalism, and economics. Its principal contribution lies in enabling heterogeneous role configurations and behavioral variability, though its validation methodologies are comparatively underdeveloped.
@@
Another significant contribution in this domain is S3 (Social-network Simulation System), which specifically focuses on social network dynamics through LLM-empowered agents [26]. S3 demonstrates particular strengths in simulating three pivotal aspects of human social behavior: emotion, attitude, and interaction behaviors. It leverages advanced prompt engineering and prompt tuning techniques to ensure authentic human-like agent behavior within social networks.
@@
Several more recent platforms have further advanced the field in terms of architectural modularity and methodological rigor. GenSim, for instance, is a general-purpose simulation engine that supports tens of thousands of LLM agents [72]. It introduces structured abstractions for social routines, long-term memory management, and adaptive behavior correction, all orchestrated by a dynamic scheduling system. Validation combines empirical replication, face validity assessments, and anomaly detection to ensure both plausibility and statistical consistency.
@@
Another emerging contribution is AgentTorch, a modular and scalable framework that unifies large-scale agent-based modeling with differentiable programming and LLM-powered behavior synthesis [15]. Instead of relying on expensive real-time prompting for each agent, AgentTorch uses a small number of archetypal LLM agents to generate compact, reusable policy representations. These are then deployed across massive populations of lightweight agents in GPU-accelerated simulations. Whereas narrative-first systems emphasize fine-grained agent individuality, AgentTorch focuses on experimental control, statistical robustness, and scientific reproducibility, making it particularly suited for large-scale hypothesis testing and parameter exploration in computational social science.
@@
SALLMA (Scalable Architecture for LLM Multi-Agent Applications) adopts a layered architecture, separating operational processes (intent formation, task execution, communication) from knowledge-level components (agent profiles, shared memory, workflows) [8]. This modular design facilitates agent specialization and integration with external tools via APIs. Its validation pipeline combines automated benchmarking with expert-in-the-loop review, aided by interactive dashboards and simulation playback for iterative refinement.
@@
SocioVerse represents a shift toward population-scale calibration [85]. It initializes agents from millions of real-world user profiles to mirror authentic distributions of attitudes, preferences, and social structures. The platform supports fine-grained tuning for scenario exploration, such as testing the effects of policy changes across diverse subpopulations. Its validation strategy emphasizes both empirical benchmarking and participatory evaluation, particularly focusing on the accurate representation of minority groups and on interactive adjustment of simulation parameters.
@@
Similarly, the LLM-AIDSim framework integrates LLMs (Llama3:8b) into ABM for influence diffusion, enabling language-level responses and deeper insights into user agent interactions [84].
@@
3.2 Architecture Patterns for Cognitive Agents
LLM-based multi-agent models generally instantiate each agent as a distinct LLM instance or session using prompt engineering. Architectures typically involve three interconnected modules: a memory system that encodes and retrieves past experiences; a reflection and summarization layer that condenses observations into mental models; and a planning component that uses these reflections to generate actions [54, 46]. These modules are often implemented as chained or nested prompts, with each step leveraging the LLM’s generative capacity to simulate cognitive processes such as recall, reasoning, and intention formation. For instance, the Humanized Agent-Based Models (h-ABM) framework, proposed by [6], outlines a modular architecture for integrating LLMs into ABMs to achieve more cognitively and emotionally realistic agents.
3.2.1 Memory Systems and Psychological Foundations
Many LLM-based agent architectures draw inspiration from cognitive psychology, particularly theories of episodic and semantic memory. For example, the memory systems in Generative Agents and AgentSociety are designed to mimic human-like recall, using vector databases or embedding-based retrieval to surface relevant past experiences. Some frameworks explicitly reference psychological models, such as the Atkinson-Shiffrin multi-store model or the concept of “theory of mind,” to inform the design of agent cognition and social reasoning [46, 21]. Reflection modules often prompt the LLM to summarize recent events, extract intentions, or infer the mental states of other agents, thus operationalizing constructs from social and developmental psychology.
Established cognitive architectures such as SOAR [43] and ACT-R [3] have provided theoretical foundations for agent design. SOAR, developed by John Laird, Allen Newell, and Paul Rosenbloom at Carnegie Mellon University [43], embodies the problem space hypothesis, suggesting that all goal-oriented behavior can be characterized as a search in a problem space defined by states, operators, and goals. This cognitive architecture divides knowledge into procedural memory (skills and know-how), semantic memory (facts about the world), and episodic memory (memories of experiences). Similarly, ACT-R (Adaptive Control of Thought-Rational) [3] posits that human cognition can be modeled through interconnected cognitive and perceptual modules that operate according to rational principles. These architectures inform the contemporary design of LLM agents by providing structured frameworks for organizing memory, decision-making, and learning.
3.2.2 Planning and Decision-Making
@@
A hybrid approach is exemplified by the LGC-MARL framework proposed by [33], which combines an LLM-based planner with graph-structured multi-agent reinforcement learning. In this architecture, the LLM is tasked with decomposing high-level instructions into executable subtasks, evaluating their coherence and feasibility, and formulating collaborative strategies among agents before final action selection. In contrast, more conventional social simulation systems rely predominantly on prompt-engineering strategies to define agent roles, goals, and reasoning pathways, typically without integrating reinforcement signals or dynamic adaptation mechanisms [36]. These approaches underscore the current design trade-off between architectural complexity and ease of interpretability in the planning components of LLM-driven simulations.
3.2.3 Orchestration and Communication
@@
Approaches to orchestration and communication vary considerably depending on the complexity and goals of the simulation. For example, platforms like Generative Agents and AgentSociety utilize orchestration layers to manage complex interactions among thousands of agents, often based on sequential flows or perception-reflection-action loops. Others, such as GenSim, integrate structured abstractions for social routines and a dynamic scheduling system. More advanced frameworks, like AgentTorch, move away from costly sequential dependencies by generating reusable policy representations from a small number of archetypal LLM agents, which are then deployed at scale on lightweight agents for GPU-accelerated simulations. SALLMA proposes a layered architecture, separating operational processes (intention formation, task execution) from knowledge-level components (agent profiles, shared memory, workflows), thereby facilitating agent specialization and integration with external tools via APIs. These platforms often support modular ”blocks,” which are reusable behavioral units that allow for systematic experimentation and increased extensibility. Integration with external APIs (e.g., web search, databases, calculators) is also crucial, as it enables agents to access additional information and tools, thereby augmenting their reasoning and decision-making capabilities [14].
@@ 3.2.4 Specialization and Behavioral Diversity
LLM-based multi-agent system (MAS) frameworks are progressively integrating sophisticated specialization protocols, allowing individual agents to adopt roles characterized by unique behavioral repertoires. This specialization is typically realized through several mechanisms: the application of tailored prompt templates that guide an agent’s responses (e.g., [80]), which can enforce specific communication styles or decision-making heuristics; the implementation of role-specific memory structures that store relevant knowledge and past interactions (e.g., [54]), enabling agents to leverage historical data pertinent to their designated function; and scenario-driven fine-tuning processes that adapt an agent’s core model to specific operational contexts (e.g., [53]), thereby optimizing their performance for particular tasks or environments. Furthermore, some cutting-edge systems are exploring methods to inject personality traits or demographic attributes into agents. This can be achieved either by conditioning prompts with specific personality descriptors (e.g., ”act as a skeptical researcher”) or by sampling attributes from real-world survey distributions (e.g., age, occupation, cultural background), thereby enhancing the realism and diversity of agent behaviors and fostering more nuanced social dynamics [77, 14]. These innovative techniques clearly enhance existing population synthesis algorithms (e.g., [12]) and enable researchers to explore complex emergent behaviors within agent populations, thereby gaining deeper insights into human-like interactions.
3.3 Current Validation Methodologies
@@ 3.3.1 Empirical Validation and Benchmarking
The validation of LLM-based social simulations remains an active and evolving area of research. Among the most rigorous efforts, frameworks like AgentSociety adopt multi-faceted validation strategies that compare simulated outputs with real-world data. Empirical benchmarking typically involves aligning simulation results with social indicators such as voter turnout, economic behavior, or survey-derived attitudinal distributions from established instruments like the General Social Survey or the World Values Survey [56]. In addition, some studies replicate classical experimental paradigms—such as public goods games or rumor propagation scenarios—within the simulation environment, and evaluate whether the emergent outcomes converge toward established empirical findings.
Longitudinal consistency is also a critical focus, particularly in simulations involving extended time horizons. Researchers analyze whether agent behavior remains stable and coherent across simulation runs, and whether emergent patterns are reproducible under minor changes to initial conditions or prompt phrasing [81]. These approaches aim to assess not only face plausibility but also the robustness and generalizability of the simulated social dynamics.
@@ 3.3.2 Human-in-the-Loop Evaluation
Calibration against real-world survey data constitutes another widely used strategy. Here, demographic or attitudinal profiles from empirical sources are used to condition LLM agents, and human evaluators subsequently assess the extent to which the simulated agents reflect the intended population characteristics [14]. Nevertheless, recent findings suggest that human raters may be influenced by non-epistemic factors—such as assertiveness, linguistic fluency, or the perceived authority of LLM outputs—leading to potential overestimation of their factual or analytical accuracy [13].